OpenStudy (anonymous):

The slant height of the cone itself is 4 inches, and the radius of the ice cream scoop on top is 0.5 inches. What is the approximate surface area of the entire ice cream cone with the scoop of ice cream?

OpenStudy (anonymous):

πr2 + πrl 3.14(0.5^2) + 3.14(0.5 * 4) Can you solve it? And if you don't have calculator then use this website calculator http://www.online-calculator.com/

OpenStudy (anonymous):

Im sorry, i dont know how to plug all the numbers into the calculator

OpenStudy (anonymous):

all right i can explain... 3.14(0.5^2) + 3.14(0.5 * 4) 3.14 * 0.25 + 3.14 * 2 Now can you solve it?

OpenStudy (anonymous):

(0.5 ? 2) im unsure what that symbol means

OpenStudy (anonymous):

In order to solve this, you need to know the equations to calculate the surface areas of a cone and a sphere. We will do this in two parts. 1) Surface area of the cone - the cone is made of a circular base so you can find the radius from this section. - the diagonal of the cone is referred to as the slant height and is often referred with 'l' Surface area of cone = pi*r*l = pi*3.5*0.75 = 2.625pi 2) Surface area of half sphere - formula for calculating surface area of sphere is 4pi*(radius)^2 - since you have a semi-sphere, your formula will become 2pi∗r2 = 1.125pi 3) Add the answers from one 1 and 2 Surface area = 3.75pi squared inches, which when rounded up will be 3.8pi squared inches.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

welcome